solve the equation:
sin(5x-π/6)=1/√2

Answer by stanbon(75887)

solve the equation:
sin(5x-π/6)=1/√2
-----
Since sin(pi/4) = 1/sqrt(2), 5x-pi/6 = pi/4
5x = pi/4 + pi/6
----
5x = 10pi/24
x = 2pi/24
x = pi/12
--------------
Cheers,
Stan H.
-------------